OpenStudy (anonymous):

FAN AND MEDAL. algebra review help ): If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. -x + 5y = 1 2x + 4y = -4 2(-5y + 1) + 4y = -4 2x + 4(-5y + 1) = -4 2x + 4(5y - 1) = -4 2(5y - 1) + 4y = -4

OpenStudy (anonymous):

1) -x + 5y = 1 2) 2x + 4y = -4 From 1): -x = 1 - 5y 3) x = 5y - 1 From 2): 2x + 4y = -4 2(5y - 1) + 4y = -4 10 y - 2 + 4y = -4 14 y = -4 + 2 14 y = -2 y = -1/7 From 3): x = 5y - 1 x = 5(-1/7) - 1 x = -5/7 - 1 x = -12/7

OpenStudy (campbell_st):

well you are making x the subject in the 1st equation add x to both sides 5y = 1 + x now subtract 1 x = 5y - 1 now substitute this into the 2nd equation.

OpenStudy (anonymous):

I substitute it into y right?

OpenStudy (anonymous):

You substitute the x = 5y - 1 into the x in 2x + 4y = -4, so that you get: 2(5y - 1) + 4y = -4

OpenStudy (anonymous):

youre awesome. thank you (:

OpenStudy (anonymous):

I'm more than happy to help! :) @campbell_st did most of the explaining though

OpenStudy (anonymous):

fanned both <3